# Complementary Probes of Warped Extra Dimension: Colliders, Gravitational Waves and Primordial Black Holes from Phase Transitions

### v1

@@ II.2 Low-energy setup: $\rho\lesssim$ TeV
In the previous setup, the SM is localized at the IR brane. This is not a viable option if the IR energy scale, $\rho$, is much below the EW scale. At the same time, localizing the SM at the UV brane is not a good alternative neither, as it does not solve the hierarchy problem at all. We tackle this issue by introducing a third (SM) brane $\mathcal{B}_{T}$ located at $z=z_{T}=1/\rho_{T}$, with $z_{0}<z_{T}<z_{1}=1/\rho$, with the goal of achieving $\rho_{T}\gtrsim 1\,\textrm{TeV}$ and $\rho_{T}\gg\rho$, and localizing the SM there Lykken and Randall (2000); Hatanaka et al. (1999); Kogan et al. (2000); Gregory et al. (2000); Kogan et al. (2001); Lee et al. (2022); Girmohanta et al. (2023). With this goal in mind, we dub $\mathcal{B}_{T}$ as TeV brane.
@@
Stabilizing the positions of the TeV and IR branes requires breaking the conformal invariance. We do it by repeating the rationale of the previous setup for both the TeV and the IR brane. We then introduce the superpotentials $W(\phi)=6k_{1}M_{5}^{2}+u_{1}k_{1}\phi^{2}$ in Region 1, and $W(\phi)=6k_{2}M_{5}^{2}+u_{2}k_{2}\phi^{2}$ in Region 2, where Region 1 and Region 2 are the two bulk regions between the branes, $z_{0}\leq z\leq z_{T}$ and $z_{T}\leq z\leq z_{1}$, respectively. On the branes, the heavy scalar field has brane potentials $\Lambda_{a}(\phi)$ analogous to those in Eq. (3), with $a=0,T,1$.
@@
In Region 1, the curvature is associated with the parameter $k_{1}$, related to the number of degrees of freedom $N_{1}$ of the sector described by the $\mathcal{B}_{T}$ brane. The location of the brane $\mathcal{B}_{T}$ with respect to the $\mathcal{B}_{0}$ brane is stabilized as $v_{0}=v_{T}(z_{0}/z_{T})^{u_{1}}$ by a heavy radion $\chi_{1}(x)$ (describing excitations $\tilde{\chi}_{1}(x)\equiv\chi_{1}(x)-\rho_{T}$ around the location of the $\mathcal{B}_{T}$ brane), which acquires vacuum expectation value $\langle\chi_{1}\rangle=\rho_{T}$. This radion is the zero mode of a 5D field with a bulk profile localized toward the $\mathcal{B}_{T}$ brane.
@@
In Region 2, the curvature and number of degrees of freedom of the dual theory associated with the $\mathcal{B}_{1}$ brane are given by $k_{2}$ and $N_{2}$. The location of the brane is stabilized by the potential of the light radion $\chi_{2}(x)$ (which similarly describes excitations $\tilde{\chi}_{2}(x)\equiv\chi_{2}(x)-\rho$ around the location of the $\mathcal{B}_{1}$ brane), and is given by $v_{T}=v_{1}(z_{T}/z_{1})^{u_{2}}$. The vacuum expectation value of this radion is $\langle\chi_{2}\rangle=\rho$. The 5D profile of the radion $\chi_{2}$ is localized toward the $\mathcal{B}_{1}$ brane.
@@
The setup thus exhibits two radion FOPTs, one at the scale $\rho_{T}\gtrsim$ TeV and one at the scale $\rho\lesssim 1\,$TeV. In this paper, we are only interested in the latter FOPT whenever we consider this setup. [^2] We postpone the full analysis of both phase transitions in the general parameter setup for future work. Therefore, for simplicity, we assume the particular case $k_{1}\simeq k_{2}=k$, (i.e. $N_{1}\simeq N_{2}=N$) and $u_{1}\simeq u_{2}=u$. In this case, the heavy radion $\chi_{1}$ is very massive Lee et al. (2022), and decouples from the low-energy theory. On the contrary, the light radion $\chi_{2}\equiv\chi$, which stabilizes the $\mathcal{B}_{1}$ brane by acquiring the vacuum expectation value $\rho$, behaves like the radion of the previous setup. The location of the two brane are fixed by the dynamical relations $v_{T}=v_{1}(z_{T}/z_{1})^{u}$ and $v_{0}=v_{1}(z_{0}/z_{1})^{u}$. Setting $\bar{v}_{0}<\bar{v}_{T}<\bar{v}_{1}$ with values of $\mathcal{O}(1)$ then naturally lead to the huge hierarchy among the Planck, TeV and IR scales, i.e. $\rho_{T}\gtrsim 1\,\textrm{TeV}$ and $\rho\lesssim 1\,$TeV.
@@
On top of the above fields, the excitations of the graviton field $h_{\mu\nu}$ propagate in the bulk as in the high-energy setup, with a massless zero mode and KK modes with masses $m_{h}\propto\rho$ and profiles localized toward the $\mathcal{B}_{1}$ brane. Their coupling with the SM fields is then further suppressed by the distance between the $\mathcal{B}_{1}$ and $\mathcal{B}_{T}$ branes $z_{T}/z_{1}$ Lee et al. (2022).
II.3 Setups summary
@@
The two setups exhibit a similar phenomenology once, in the low-energy setup with the hierarchy $z_{1}/z_{T}\gg 1$, the heavy radion is decoupled and the heavy-radion dynamics does not impact the light-radion FOPT. Indeed, the KK excitations only appear for the fields propagating in the bulk and, since the SM particles are bound in a brane, only the bulk field $\phi$ and the metric provide 5D resonances. However, $\phi$ does not provide an independent degree of freedom, so that the gravitons and the radion of the IR brane drive the BSM signatures of the model. Their phenomenology depends on the following free parameters: $N$, $\rho$, $\lambda_{1}$, $u$ and $v_{a}$. As Eqs. (4) and (6) show, $N$, $\rho$ and $\lambda_{1}$ control the energy gap $E_{0}$, which the FOPT inverse duration parameter $H_{*}/\beta$ depends on, while $u$ and $v_{a}$ control the (dimensionless) IR-brane radion mass $\hat{m}_{\chi}$. Our analysis can then use $N$, $\rho$, $\beta/H_{*}$ and $\hat{m}_{\chi}$ as proxies of the aforementioned free parameters.
III Strong first-order phase transition
@@
The phase transition can start at temperatures $T<T_{c}$ where the critical temperature, $T_{c}$, is defined as $F_{c}(T_{c})=F_{d}(T_{d})$, which leads from Eq. (7) to
@@
with boundary conditions $\chi(0)=\chi_{0}$ and $d\chi/d\sigma=0$. The corresponding temperature is obtained by equating the kinetic energy of the radion at the origin, $\chi=0$, with the thermal energy, so that the barrier between the two vacua can be overcome  [^5] This approach is an excellent approximation of the full bounce solution, including both the radion and the degree of freedom associated to the position of the AdS-Swarzschild horizon. For details see e.g. Ref. Megías et al. (2020)..
By definition, the nucleation temperature $T_{n}$ is reached when there is one bubble produced by causal Hubble volume. At this temperature, the action has value $S_{4}(T_{n})=S_{n}$, where $S_{n}$ is computed in App. A. To determine $T_{n}$, we then solve numerically Eq. (11) and check that the condition $S_{4}(T_{n})=S_{n}$ is met. In App. B we compare this approach with the thick-wall approximation, which is supposed to be reliable for supercooled FOPT. The comparison shows that the thick-wall approximation does not describe accurately the numerical results in all our benchmark scenarios, so that, in order to simplify the numerical analysis, we implement semi-analytical (thick-wall inspired) $\ae$ for the different temperatures that fit the numerical result in all the considered cases.
@@
In our thick-wall inspired expressions, we limit ourselves to $\beta/H_{\ast}\lesssim 10$, for which the formation of PBH is relevant, as we will see in the next section  [^6] For moderately strong FOPT, we do not need to distinguish between the Hubble parameter at the nucleation temperature $H_{n}$, and the Hubble parameter during percolation $H_{\ast}$, as their values are very similar Caprini et al. (2020). In fact, in our model and for the considered values of $\beta/H_{*}$ we can see from the middle panel of Fig. 1 that both temperatures are very close to each other.. As for the energy balance, we quantify it by the parameter $\alpha(T_{n})$ given by
@@
Table 2: Benchmark points for the FOPT phenomenology analysis. In the top block, the input values. In the next-to-top block, the outputs about the FOPT quantities detailed in Sec. III. In the third block, the outputs concerning the PBH quantities defined in Sec. IV. In the bottom block, the outputs on the SGWB predictions discussed in Sec. V. For a given set of $\{N,\rho\}$ input, the value of $\lambda_{1}$ is chosen to maximize the PBH abundance experimentally allowed, as explained in Sec. IV.
IV Primordial black hole formation
A FOPT can lead to formation of PBHs in different ways Liu et al. (2022); Kawana et al. (2023); Gouttenoire and Volansky (2024); Lewicki et al. (2023). In particular, due to the stochastic nature of nucleation, there is a probability that some regions remain for a longer time in the false vacuum while the space around them gets filled by true vacuum. Reheating in the true vacuum leads to radiation, so that its energy density $\rho_{\rm true}$ dilutes, while the energy density in the false vacuum is constant $\rho_{\text{false}}\simeq V_{\Lambda}$ if the FOPT transition is extremely supercooled. Thus regions in the false vacuum become relatively denser as quantified by $\delta\equiv\rho_{\text{false}}/\rho_{\text{true}}-1$, with the effect being more pronounced if the time gap between nucleation and percolation is long, i.e. the FOPT is slow or, equivalently, $\beta/H$ is small. If a region in the false vacuum (approximated as roughly-spherical) has radius larger than roughly the Hubble parameter, it forms a PBH when the density contrast exceeds the critical value $\delta_{\text{c}}\approx 0.45$ Carr et al. (2021). This means that PBHs can form due to the collapse of overdense regions during a supercooled FOPT Gouttenoire and Volansky (2024); Conaci et al. (2024).
In the limit of $\alpha\gg 1$, the probability that a Hubble patch collapses into a PBH, $\mathcal{P}_{\rm coll}$, can be approximated by the analytic formula Gouttenoire and Volansky (2024); Conaci et al. (2024)
@@ IV Primordial black hole formation
where $M_{\odot}\simeq 2\times 10^{30}\,\text{kg}$ is the solar mass Gouttenoire and Volansky (2024). Their abundance, expressed as the fraction of DM in the form of PBH today, is given by
@@
Their spin can be estimated using the peak theory formalism and is parameterized by the dimensionless Kerr parameter $a_{\ast}$ Banerjee and Dey (2024); Conaci et al. (2024); Arteaga et al. (2024):
@@
In the left panel of Fig. 2 we summarize the PBH experimental bounds and future probes. The gray area corresponds to the parameter region $\{f_{\rm PBH},M_{\rm PBH}\}$ experimentally ruled out in scenarios where the PBH mass distribution is extremely peaked, say monochromatic, which is what we expect in our radion FOPT. The gray area includes the following bounds (see Refs. Green and Kavanagh (2021); Saha and Laha (2022); Laha (2019); Ray et al. (2021) for details):
  • Hawking evaporation of PBH is relevant at masses $5\times 10^{-24}M_{\odot}\lesssim M_{\rm PBH}\lesssim 10^{-16}M_{\odot}$ and implies constraints using data from CMB Clark et al. (2017), EDGES Mittal et al. (2022), IntegralLaha et al. (2020); Berteaud et al. (2022), VoyagerBoudaud and Cirelli (2019), 511 keV DeRocco and Graham (2019) and EGRB Carr et al. (2010).
  • Micro-lensing observations from HSC Niikura et al. (2019a), EROS Tisserand et al. (2007), IcarusOguri et al. (2018), including a hint for PBH in OGLE Niikura et al. (2019b), are relevant for masses $5\times 10^{-11}M_{\odot}\lesssim M_{\rm PBH}\lesssim 10^{4}M_{\odot}$.
  • LVK detections binds the properties of the stellar-mass black population, imposing a constraint on PBHs with $1\,M_{\odot}\lesssim M_{\rm PBH}\lesssim 100\,M_{\odot}$ Franciolini et al. (2022a); Kavanagh et al. (2018); Hall et al. (2020); Wong et al. (2021); Hütsi et al. (2021); De Luca et al. (2021a); Franciolini et al. (2022b).
  • Heavy PBHs are expected to accrete, leading to signatures in the CMB. The null observation of this effect binds the region $M_{\rm PBH}\gtrsim 10\,M_{\odot}$ Serpico et al. (2020); Piga et al. (2022).
The colored areas instead corresponds to the PBH parameter reach achivable at forthcoming experiments. Micro-lensing future sensitivity of the NGRST DeRocco et al. (2024) is depicted in red. ET and LISA sensitivities to BPH mergers are in green and blue, respectively Pujolas et al. (2021) (see also Refs. De Luca et al. (2021b); Franciolini et al. (2022c); Martinelli et al. (2022); Franciolini et al. (2023); Marcoccia et al. (2024); Branchesi et al. (2023); Marriott-Best et al. (2024)). Therefore, the PBH mass window able to explain the entire DM abundance is
@@
By looking at Eqs. (22), (23) and (24), one can notice that $f_{\rm PBH}$ only depends on $\beta/H_{\ast}$ once $M_{\text{PBH}}$ (or equivalently, $T_{R}$) is fixed. Moreover, $M_{\rm PBH}$ (or equivalently $T_{R}$) is a function of $\rho$, $N$ and $\beta/H_{*}$ via Eqs. (16) and (48). This allows us to map, for a fixed $N$, every point of the PBH space parameter $\{f_{\rm PBH},M_{\rm PBH}\}$ to a point of the plane $\{f_{\rm PBH},\rho\}$ or to the plane $\{\rho,\beta/H_{\ast}\}$. Using these mappings, in Fig. 2 we recast the aforementioned constraints and future sensitivities given in the plane $\{f_{\rm PBH},M_{\rm PBH}\}$ (left panel) in terms of other parameters (middle and right panels) for $N=10,20,50$ (dashed, solid, dotted borders). As the central panel shows, in our (high-energy) 5D setup the PBH mass window in Eq. (26) corresponds to $\rho\in[33.2,2.40\times 10^{4}]\,\textrm{TeV}$ for $N=10$, $\rho\in[16.6,1.20\times 10^{4}]\,\textrm{TeV}$ for $N=20$, and $\rho\in[6.6,4.8\times 10^{3}]\,\textrm{TeV}$ for $N=50$, which arises for $15\,\textrm{TeV}\lesssim T_{R}\lesssim 10\,{\rm PeV}$. The comparison of the middle and right panels moreover highlights the extreme sensitivity of the parameter $f_{{\rm PBH}}$ to $\beta/H_{\ast}$ due to the exponential dependence in Eq. (22). This sensitivity introduces a tuning $\Delta\sim 10^{-4}$ in the model if one aims at achieving $f_{\rm PBH}\sim 1$ and defines the tuning parameter $\Delta$ as
$\Delta=\textrm{max}_{p}\Delta_{p}\simeq\Delta_{\lambda_{1}}=\left|\frac{d\log f_{\rm PBH}}{d\log\lambda_{1}}\right|\qquad\textrm{with}\qquad p=\rho,N,\lambda_{1}\qquad\textrm{and}\qquad\Delta_{p}=\left|\frac{d\log f_{\rm PBH}}{d\log p}\right|\,,$ (27)
where the dependence of $f_{\rm PBH}$ on $\lambda_{1}$ is obtained from Eqs. (20), (22) and (24). The fine-tuning in our model ($\gtrsim 10^{-4}$) turns out to be several orders of magnitude less severe than that required to achieve $f_{\rm PBH}\simeq 1$ in typical single-field inflationary models Cole et al. (2023). The border line of the gray area in Fig. 2 defines the maximal fraction of PBH, $f_{\rm PBH}^{\rm max}$ (left and middle panels), or the minimal duration of FOPT $(\beta/H_{\ast})_{\rm min}$ (right panel), which is experimentally allowed for a monochromatic population of mass PBH $M_{\rm PBH}$. The black dots along this line correspond to the benchmark points presented in Table 2, in which we also quote the values of the PBH quantities obtained with the above expressions. Increasing $-\lambda_{1}$ of a benchmark point would correspond to moving on the vertical dashed line below the black dot of that given benchmark.
@@
Figure 3: Left panel: $\rho$ as a function of $M_{{\rm PBH}}/M_{\odot}$ for $N=10,20,50$ (dashed, solid, dotted line). When varying $M_{{\rm PBH}}$, $f_{\rm PBH}$ is adjusted to the maximal value that is experimentally allowed for that value of $M_{{\rm PBH}}$ (cf. border line of the gray area in Fig. 2). Middle panel: Contour lines of $f_{{\rm PBH}}$ in the plane $\{\rho,N\}$. Right panel: Contour plot of the BH spin, $S_{{\rm PBH}}$, as given by Eq. (25), in the plane $\{\rho,N\}$. In the middle and right panels, we have considered $\beta/H_{\ast}=(\beta/H_{\ast})_{\rm min}$. The colors in the middle and right panels correspond to density plots to better appreciate the regions with higher (reddish) and lower (bluish) values of the observables.
We show in the middle panel of Fig. 3 contour lines of $f_{\rm PBH}$ in the plane $\{\rho,N\}$ for $\beta/H_{\ast}=(\beta/H_{\ast})_{\rm min}$, thus avoiding any forbidden region from the present experiments, but that can be probed by future experiments NGRST, LISA, and ET.

### v2

@@ 2.2 Low-energy setup: $\rho\lesssim$ TeV
In the previous setup, the SM is localized at the IR brane. This is not a viable option if the IR energy scale, $\rho$, is much below the EW scale. At the same time, localizing the SM at the UV brane is not a good alternative neither, as it does not solve the hierarchy problem at all. We tackle this issue by introducing a third (SM) brane $\mathcal{B}_{T}$ located at $z=z_{T}=1/\rho_{T}$, with $z_{0}<z_{T}<z_{1}=1/\rho$, with the goal of achieving $\rho_{T}\gtrsim 1\,\textrm{TeV}$ and $\rho_{T}\gg\rho$, and localizing the SM there Lykken:1999nb ; Hatanaka:1999ac ; Kogan:1999wc ; Gregory:2000jc ; Kogan:2000xc ; Agashe:2016rle ; Agashe:2016kfr ; Lee:2021wau ; Girmohanta:2023sjv . With this goal in mind, we dub $\mathcal{B}_{T}$ as TeV brane.
@@
Stabilizing the positions of the TeV and IR branes requires breaking the conformal invariance. We do it by repeating the rationale of the previous setup for both the TeV and the IR brane. We then introduce the superpotentials $W(\phi)=6k_{1}M_{5}^{2}+u_{1}k_{1}\phi^{2}$ in Region 1, and $W(\phi)=6k_{2}M_{5}^{2}+u_{2}k_{2}\phi^{2}$ in Region 2, where Region 1 and Region 2 are the two bulk regions between the branes, $z_{0}\leq z\leq z_{T}$ and $z_{T}\leq z\leq z_{1}$, respectively. On the branes, the stabilizing scalar field has brane potentials $\Lambda_{a}(\phi)$ analogous to those in Eq. (3), with $a=0,T,1$.
@@
In Region 1, the curvature is associated with the parameter $k_{1}$, related to the number of degrees of freedom $N_{1}$ of the sector described by the $\mathcal{B}_{T}$ brane. The location of the brane $\mathcal{B}_{T}$ with respect to the $\mathcal{B}_{0}$ brane is stabilized as $v_{0}=v_{T}(z_{0}/z_{T})^{u_{1}}$ by a heavy radion $\chi_{1}(x)$ (describing excitations $\tilde{\chi}_{1}(x)\equiv\chi_{1}(x)-\rho_{T}$ around the location of the $\mathcal{B}_{T}$ brane, and with a mass ($m_{\chi_{1}}\sim u_{1}\rho_{T}$) parametrically controlled by the parameter $\rho_{T}$), which acquires vacuum expectation value $\langle\chi_{1}\rangle=\rho_{T}$. This radion is the zero mode of a 5D field with a bulk profile localized toward the $\mathcal{B}_{T}$ brane.
@@
In Region 2, the curvature and number of degrees of freedom of the dual theory associated with the $\mathcal{B}_{1}$ brane are given by $k_{2}$ and $N_{2}$. The location of the brane is stabilized by the potential of the light radion $\chi_{2}(x)$ (which similarly describes excitations $\tilde{\chi}_{2}(x)\equiv\chi_{2}(x)-\rho$ around the location of the $\mathcal{B}_{1}$ brane, with a mass ($m_{\chi_{2}}\sim u_{2}\rho$) parametrically controlled by the parameter $\rho$), and is given by $v_{T}=v_{1}(z_{T}/z_{1})^{u_{2}}$. The vacuum expectation value of this radion is $\langle\chi_{2}\rangle=\rho$. The 5D profile of the radion $\chi_{2}$ is localized toward the $\mathcal{B}_{1}$ brane.
@@
The setup thus exhibits two radion FOPTs, one at the scale $\rho_{T}\gtrsim$ TeV and one at the scale $\rho\lesssim 1\,$TeV. In this paper, we are only interested in the latter FOPT whenever we consider this setup. [^2] We postpone the full analysis of both phase transitions in the general parameter setup for future work. Therefore, for simplicity, we assume the particular case $k_{1}\simeq k_{2}=k$, (i.e. $N_{1}\simeq N_{2}=N$) and $u_{1}\simeq u_{2}=u$. In this case, the heavy radion $\chi_{1}$ is very massive (as $m_{\chi_{1}}^{2}\propto\frac{1}{k_{1}-k_{2}}$ee:2021wau , and decouples from the low-energy theory. On the contrary, the light radion $\chi_{2}\equiv\chi$, which stabilizes the $\mathcal{B}_{1}$ brane by acquiring the vacuum expectation value $\rho$, behaves like the radion of the previous setup. The location of the two branes are fixed by the dynamical relations $v_{T}=v_{1}(z_{T}/z_{1})^{u}$ and $v_{0}=v_{1}(z_{0}/z_{1})^{u}$. Setting $\bar{v}_{0}<\bar{v}_{T}<\bar{v}_{1}$ with values of $\mathcal{O}(1)$ then naturally lead to the huge hierarchy among the Planck, TeV and IR scales, i.e. $\rho_{T}\gtrsim 1\,\textrm{TeV}$ and $\rho\lesssim 1\,$TeV.
@@
On top of the above fields, the excitations of the graviton field $h_{\mu\nu}$ propagate in the bulk as in the high-energy setup, with a massless zero mode and KK modes with masses $m_{h}\propto\rho$ and profiles localized toward the $\mathcal{B}_{1}$ brane. Their coupling with the SM fields is then further suppressed by the distance between the $\mathcal{B}_{1}$ and $\mathcal{B}_{T}$ branes $z_{T}/z_{1}$ Lee:2021wau .
2.3 Setups summary
@@
The two setups exhibit a similar phenomenology once, in the low-energy setup with the hierarchy $z_{1}/z_{T}\gg 1$, the heavy radion is decoupled and the heavy-radion dynamics does not impact the light-radion FOPT. Indeed, the KK excitations only appear for the fields propagating in the bulk and, since the SM particles are bound in a brane, only the bulk field $\phi$ and the metric provide 5D resonances. However, $\phi$ does not provide an independent degree of freedom, so that the gravitons and the radion of the IR brane drive the BSM signatures of the model. Their phenomenology depends on the following free parameters: $N$, $\rho$, $\lambda_{1}$, $u$ and $v_{a}$. As Eqs. (4) and (6) show, $N$, $\rho$ and $\lambda_{1}$ control the energy gap $E_{0}$, which the FOPT duration parameter $H_{*}/\beta$ depends on, while $u$ and $v_{a}$ control the (dimensionless) IR-brane radion mass $\hat{m}_{\chi}$. Our analysis can then use $N$, $\rho$, $\beta/H_{*}$ and $\hat{m}_{\chi}$ as proxies of the aforementioned free parameters.
3 Strong first-order phase transition
@@
The phase transition can start at temperatures $T<T_{c}$ where the critical temperature, $T_{c}$, is defined as $F_{c}(T_{c})=F_{d}(T_{c})$, which leads from Eq. (7) to
@@
with boundary conditions $\chi(0)=\chi_{0}$ and $d\chi/d\sigma=0$. The corresponding temperature is obtained by equating the kinetic energy of the radion at the origin, $\chi=0$, with the thermal energy, so that the barrier between the two vacua can be overcome  [^5] This approach is an excellent approximation of the full bounce solution, including both the radion and the degree of freedom associated to the position of the AdS-Swarzschild horizon. For details see e.g. Ref. Megias:2020vek ..
By definition, the nucleation temperature $T_{n}$ is reached when there is one bubble produced per causal Hubble volume. At this temperature, the action has value $S_{4}(T_{n})=S_{n}$, where $S_{n}$ is computed in App. A. To determine $T_{n}$, we then solve numerically Eq. (11) and check that the condition $S_{4}(T_{n})=S_{n}$ is met. In App. B we compare this approach with the thick-wall approximation, which is supposed to be reliable for supercooled FOPT. The comparison shows that the thick-wall approximation does not describe accurately the numerical results in all our benchmark scenarios, so that, in order to simplify the numerical analysis, we implement semi-analytical (thick-wall inspired) $\ae$ for the different temperatures that fit the numerical result in all the considered cases.
@@
In our thick-wall inspired expressions, we limit ourselves to $\beta/H_{\ast}\lesssim 10$, for which the formation of PBH is relevant, as we will see in the next section  [^6] For moderately strong FOPT, we do not need to distinguish between the Hubble parameter at the nucleation temperature $H_{n}$, and the Hubble parameter during percolation $H_{\ast}$, as their values are very similar Caprini:2019egz . In fact, in our model and for the considered values of $\beta/H_{*}$ we can see from the middle panel of Fig. 1 that both temperatures are very close to each other.. As for the energy balance, we quantify it by the parameter $\alpha(T_{n})$ given by
@@
Table 2: Benchmark points for the FOPT phenomenology analysis. In the top block, the input values. In the next-to-top block, the outputs about the FOPT quantities detailed in Sec. 3. In the third block, the outputs concerning the PBH quantities defined in Sec. 4. In the bottom block, the outputs on the SGWB predictions discussed in Sec. 5. For a given set of $\{N,\rho\}$ input, the value of $\lambda_{1}$ is chosen to maximize the PBH abundance experimentally allowed, as explained in Sec. 4.
4 Primordial black hole formation
A FOPT can lead to formation of PBHs in different ways Liu:2021svg ; Kawana:2022olo ; Gouttenoire:2023naa ; Lewicki:2023ioy . In particular, due to the stochastic nature of nucleation, there is a probability that some regions remain for a longer time in the false vacuum while the space around them gets filled by true vacuum. Reheating in the true vacuum leads to radiation, so that its energy density $\rho_{\rm true}$ dilutes, while the energy density in the false vacuum is constant $\rho_{\text{false}}\simeq V_{\Lambda}$ if the FOPT transition is extremely supercooled. Thus regions in the false vacuum become relatively denser as quantified by $\delta\equiv\rho_{\text{false}}/\rho_{\text{true}}-1$, with the effect being more pronounced if the time gap between nucleation and percolation is long, i.e. the FOPT is slow or, equivalently, $\beta/H$ is small. If a region in the false vacuum (approximated as roughly-spherical) has radius larger than roughly the inverse Hubble parameter, it forms a PBH when the density contrast exceeds the critical value $\delta_{\text{c}}\approx 0.45$ 2002.12778 . This means that PBHs can form due to the collapse of overdense regions during a supercooled FOPT Gouttenoire:2023naa ; Conaci:2024tlc .
In the limit of $\alpha\gg 1$, the probability that a Hubble patch collapses into a PBH, $\mathcal{P}_{\rm coll}$, can be approximated by the analytic formula Gouttenoire:2023naa ; Conaci:2024tlc
@@ 4 Primordial black hole formation
where $M_{\odot}\simeq 2\times 10^{30}\,\text{kg}$ is the solar mass Gouttenoire:2023naa . Their abundance, expressed as the fraction of DM in the form of PBH today, is given by
@@
Their spin can be estimated using the peak theory formalism and is parameterized by the dimensionless Kerr parameter $a_{\ast}$ Banerjee:2023qya ; Conaci:2024tlc ; Arteaga:2024vde :
@@
In the left panel of Fig. 2 we summarize the PBH experimental bounds and future probes. The gray area corresponds to the parameter region $\{f_{\rm PBH},M_{\rm PBH}\}$ experimentally ruled out in scenarios where the PBH mass distribution is extremely peaked, say monochromatic, which is what we expect in our radion FOPT. The gray area includes the following bounds (see Refs. Green:2020jor ; Saha:2021pqf ; Laha:2019ssq ; Ray:2021mxu for details):
  • Hawking evaporation of PBH is relevant at masses $5\times 10^{-24}M_{\odot}\lesssim M_{\rm PBH}\lesssim 10^{-16}M_{\odot}$ and implies constraints using data from CMB Clark:2016nst , EDGES Mittal:2021egv , IntegralLaha:2020ivk ; Berteaud:2022tws , VoyagerBoudaud:2018hqb , 511 keV DeRocco:2019fjq and EGRB Carr:2009jm .
  • Micro-lensing observations from HSC Niikura:2017zjd , EROS EROS-2:2006ryy , IcarusOguri:2017ock , including a hint for PBH in OGLE Niikura:2019kqi , are relevant for masses $5\times 10^{-11}M_{\odot}\lesssim M_{\rm PBH}\lesssim 10^{4}M_{\odot}$.
  • LVK detections binds the properties of the stellar-mass black population, imposing a constraint on PBHs with $1\,M_{\odot}\lesssim M_{\rm PBH}\lesssim 100\,M_{\odot}$ Franciolini:2022tfm ; Kavanagh:2018ggo ; Hall:2020daa ; Wong:2020yig ; Hutsi:2020sol ; DeLuca:2021wjr ; Franciolini:2021tla .
  • Heavy PBHs are expected to accrete, leading to signatures in the CMB. The null observation of this effect binds the region $M_{\rm PBH}\gtrsim 10\,M_{\odot}$ Serpico:2020ehh ; Piga:2022ysp .
The colored areas instead correspond to the PBH parameter reach achivable at forthcoming experiments. Micro-lensing future sensitivity of the NGRST DeRocco:2023gde is depicted in red. ET and LISA sensitivities to BPH mergers are in green and blue, respectively Pujolas:2021yaw (see also Refs. DeLuca:2021hde ; Franciolini:2022htd ; Martinelli:2022elq ; Franciolini:2023opt ; Marcoccia:2023khb ; Branchesi:2023mws ; Marriott-Best:2024anh ). Therefore, the PBH mass window able to explain the entire DM abundance is
@@
By looking at Eqs. (22), (23) and (24), one can notice that $f_{\rm PBH}$ only depends on $\beta/H_{\ast}$ once $M_{\text{PBH}}$ (or equivalently, $T_{R}$) is fixed. Moreover, $M_{\rm PBH}$ (or equivalently $T_{R}$) is a function of $\rho$, $N$ and $\beta/H_{*}$ via Eqs. (16) and (48). This allows us to map, for a fixed $N$, every point of the PBH space parameter $\{f_{\rm PBH},M_{\rm PBH}\}$ to a point of the plane $\{f_{\rm PBH},\rho\}$ or to the plane $\{\rho,\beta/H_{\ast}\}$. Using these mappings, in Fig. 2 we recast the aforementioned constraints and future sensitivities given in the plane $\{f_{\rm PBH},M_{\rm PBH}\}$ (left panel) in terms of other parameters (middle and right panels) for $N=10,20,50$ (dashed, solid, dotted borders). As the central panel shows, in our (high-energy) 5D setup the PBH mass window in Eq. (26) corresponds to $\rho\in[33.2,2.40\times 10^{4}]\,\textrm{TeV}$ for $N=10$, $\rho\in[16.6,1.20\times 10^{4}]\,\textrm{TeV}$ for $N=20$, and $\rho\in[6.6,4.8\times 10^{3}]\,\textrm{TeV}$ for $N=50$, which arises for $15\,\textrm{TeV}\lesssim T_{R}\lesssim 10\,{\rm PeV}$. The comparison of the middle and right panels moreover highlights the extreme sensitivity of the parameter $f_{{\rm PBH}}$ to $\beta/H_{\ast}$ due to the exponential dependence in Eq. (22).
This sensitivity introduces a tuning $\Delta\sim 10^{-4}$ in the model if one aims at achieving $f_{\rm PBH}\sim 1$ and defines the tuning parameter $\Delta$ as
$\Delta=\textrm{max}_{p}\Delta_{p}\simeq\Delta_{\lambda_{1}}=\left|\frac{d\log f_{\rm PBH}}{d\log\lambda_{1}}\right|\quad\textrm{with}\quad p=\rho,N,\lambda_{1}\quad\textrm{and}\quad\Delta_{p}=\left|\frac{d\log f_{\rm PBH}}{d\log p}\right|\,,$ (27)
where the dependence of $f_{\rm PBH}$ on $\lambda_{1}$ is obtained from Eqs. (20), (22) and (24). The fine-tuning in our model ($\gtrsim 10^{-4}$) turns out to be several orders of magnitude less severe than that required to achieve $f_{\rm PBH}\simeq 1$ in typical single-field inflationary models Cole:2023wyx . The border line of the gray area in Fig. 2 defines the maximal fraction of PBH, $f_{\rm PBH}^{\rm max}$ (left and middle panels), or the minimal duration of FOPT $(\beta/H_{\ast})_{\rm min}$ (right panel), which is experimentally allowed for a monochromatic population of PBH with mass $M_{\rm PBH}$. The black dots along this line correspond to the benchmark points presented in Table 2, in which we also quote the values of the PBH quantities obtained with the above expressions. Increasing $-\lambda_{1}$ of a benchmark point would correspond to moving on the vertical dashed line below the black dot of that given benchmark.
@@
Figure 3: Upper left panel: $\rho$ as a function of $M_{{\rm PBH}}/M_{\odot}$ for $N=10,20,50$ (dashed, solid, dotted line). When varying $M_{{\rm PBH}}$, $f_{\rm PBH}$ is adjusted to the maximal value that is experimentally allowed for that value of $M_{{\rm PBH}}$ (cf. border line of the gray area in Fig. 2). Upper right panel: Contour plot of the BH spin, $S_{{\rm PBH}}$, as given by Eq. (25), in the plane $\{\rho,N\}$. Lower panel: As for the upper right panel but for $f_{{\rm PBH}}$. The parameter reach for PBH detection by the LISA, ET, and NGRST experiments is included. In the upper right and lower panels, we have considered $\beta/H_{\ast}=(\beta/H_{\ast})_{\rm min}$.